OpenStudy (anonymous):

Can these functions be composed together using fog?: f(x)=x^3 - 3x^2 + 2x - 5, g(x)= rad(x) So far I have: rad(x^3) - 3 rad(x^2) + 2 rad(x) - 5...

OpenStudy (amistre64):

if you "want" to combine these equation into f(g(x)) it is possible. f(g(x)) = (rad(x))^3 - 3(rad(x))^2 + 2(rad(x)) - 5 what do you mean by rad(x)? is it a square root radical?

OpenStudy (anonymous):

Thank you for your help...I mean Radical x. I am getting the same thing... can it be simplified further?

OpenStudy (amistre64):

if the radical represents a square root then yes, it can be simplified; but for clarity, let use 2root(x) to mean square root... ok?

OpenStudy (amistre64):

(2root(x))^3 = x(2root(x)) -3(2root(x))^2 = -3x those are tho only "simplifications to that part...

OpenStudy (amistre64):

f(g(x)) = x(2root(x)) - 3x + 2(2root(x)) - 5 2root(x) (x+2) -3x -5 ?

OpenStudy (anonymous):

That looks good! it matches this: (fog)(x)= (x+2) 2root(x) -3x - 5. Thank you!